OpenStudy (anonymous):

can someone please help me: what is x^2 + 14x + 48 factored???!

OpenStudy (whpalmer4):

To factor \(x^2+14x+48\) we'll need to find two binomials to multiply together so that \[(x+a)(x+b) = x^2 + 14x + 48\] Let's expand \((x+a)(x+b)\) to get\[(x+a)(x+b) = x^2 + ax + bx + ab = x^2 + (a+b)x + ab\] Look at that for a minute. That means if we can think of two numbers, \(a, b\), such that \(a+b = 14\) and \(a * b = 48\) we can just plug them in and be done! What are all the pairs of numbers that multiply together to give you 48?